write the equation of a line whose slope is -2 and passes through (-1,6). write the answer in slope intercept form.
This one I have no clue about.

Answer by checkley71(8403)

Y=mX+b WHERE Y & X ARE COORDINATES OF A POINT ON THIS LINE, m=SLOPE & b=Y INTERCEPT.
6=-2*-1+b
6=2+b
b=6-2
b=4 FOR THE Y INTERCEPT.
THUS THE LINE EQUATION IS:
Y=-2X+4